Grade 12Algebra

Find:
(2/5) ÷ (½)
Find:
(2/5) ÷ (½)
Find:
(2/5) ÷ (½)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (2/5) × reciprocal of 1⁄2
= (2/5) × (2/1)
= (2 × 2) / (5 × 1)
= 4/5

Thanks